ADVANCEMENTS IN PHOTOMIXING AND PHOTCONDUCTIVE SWITCHING FOR THZ SPEC-TROSCOPY

\underline{E.R. BROWN}, Physical Domains, LLC, P.O. Box 1574, La Canada, CA 91012; JOSEPH R. DE-MERS, EMCORE Corp., 2015 W. Chestnut St., Alhambra, CA 91803.

Ultrafast photoconductive devices have been an important breakthrough in THz technology during the past two decades. Photoconductive switches have become the workhorse in moderate-resolution time-domain systems, and photomixers have been widely implemented in high-resolution spectrometers of various types. The primary photoconductive material has been low-temperature-grown GaAs. More recently, this has been eclipsed by ErAs-GaAs: a nanocomposite consisting of ErAs nanoparticles embedded in a GaAs matrix. ErAs-GaAs has produced very useful THz output power levels when pumped by low-cost lasers operating around 780 nm. Semiconducting DFB lasers have produced cw photomixer output levels between roughly 10 and 1 microwatt between 0.1 and 1.0 THz, respectively. Frequency-doubled fiber mode-locked lasers having an average output power of 20 mW have produced average THz output powers approaching 1 mW and peak power exceeding 1 W. This photomixer performance has been utilized in the first commercial THz photomixing spectrometer manufactured by EMCORE Corp., which has already been demonstrated on a variety of interesting materials including polar vapors, solid explosives, polysaccharides, nucleic acids, and non-centrosymmetric crystals. Results from each of these types will be presented along with resolution and acquisition-time performance metrics for the latest photomixing system.